OpenStudy (anonymous):

Help with extended proportion? x : y : z = 2 : 3 : 7 ; 4x + 5y - 3z = 6 x/2 = y/3 = z/7 (=k) x/2 = k ; x = 2k y/3 = k ; y = 3k z/7 = k ; z = 7k 2k + 3k + 7k = 6 12k = 6 k= ? Did I do something wrong?

OpenStudy (kc_kennylau):

Yes you did x+y+z=6 instead of 4x+5y-3z=6 i.e. incorrect substitution

OpenStudy (anonymous):

I have to find the value of k before doing 2*k = x ; k*3 = y ; k*7 = z then I can use the values of those to substitute in the equation. I do not have k yet.

OpenStudy (kc_kennylau):

You substituted x=2k, y=3k, z=7k into x+y+z=6 instead of 4x+5y-3z=6

OpenStudy (anonymous):

Here is an example of what my professor has done. x : y : z : t = 1 : 2 : 3 : 4 x+y+z+t = 190. x/1 = k, x= k y/2 = k, y = 2k z/3 = k; z=3k t/4 = k; t=4k k + 2k + 3k + 4k = 190 10k = 190 k = 19 Now k*1 = 19 so x = 19 2*19 = 38 -> y = 38 3*19 = 57 so z = 57 4*19 = 76 so t = 76 that is why : 19 + 38 + 57 + 76 = 190.

OpenStudy (kc_kennylau):

Because the equation was x+y+z+t=190

OpenStudy (kc_kennylau):

Now the equation is 4x+5y-3z=6

OpenStudy (kc_kennylau):

He was substituting x=k,y=2k,z=3k,t=4k into x+y+z+t=190

OpenStudy (anonymous):

Yes but the question I am doing I HAVE NOT FOUND K yet therefore I DID NOT substitute ANYTHING yet.

OpenStudy (anonymous):

how would I find k? do I have to multiply 2k by 4 or?

OpenStudy (kc_kennylau):

Then explain to me how you got 2k + 3k + 7k = 6

OpenStudy (anonymous):

I probably have forgot to multiply them?

OpenStudy (kc_kennylau):

Exactly

OpenStudy (anonymous):

so this means 2k * 4 3k*5 and 7k * 3??

OpenStudy (kc_kennylau):

Yes

OpenStudy (anonymous):

Ohhh now i get it! thank you! let me see if I can get the correct answer now

OpenStudy (kc_kennylau):

No problem :)

OpenStudy (kc_kennylau):

x=6,y=9,z=21

OpenStudy (anonymous):

I got the answer! :D

OpenStudy (kc_kennylau):

Glad you got the answer

OpenStudy (anonymous):

Yep i did another question the same and got it too :D